I need help with this problem. I'm out of my depth with this. Thank you kindly.

Morton Mercado, the CFO of Kanton Company, carefully developed the estimates of the firm's total funds requirements for the coming year. These are shown in the following table:

Month

Total Funds

Month

Total Funds

January

$1,000,000

July

$6,000,000

February

$1,000,000

August

$5,000,000

March

$2,000,000

September

$5,000,000

April

$3,000,000

October

$4,000,000

May

$5,000,000

November

$2,000,000

June

$7,000,000

December

$1,000,000

In addition, Morton expects short-term financing costs of about 10% and long-term financing costs of about 14% during that period. He developed the three possible financing strategies that follow:

Strategy 1 - Aggressive: Finance seasonal needs with short-term finds and permanent needs with long-term funds.

Strategy 2 - Conservative: Finance an amount equal to the peak need with long-term funds and use short-term funds only in an emergency.

Strategy 3 - Tradeoff: Finance $3,000,000 with long-term funds and finance the remaining funds requirements with short-term funds.

Using data on the firm's total funds requirements, Morton estimated the average annual short-term and long-term financing requirements for each strategy in the coming year, as shown in the following table.

AVERAGE

ANNUAL

FINANCING

Type of Financing

Strategy 1 Aggressive

Strategy 2 Conservative

Strategy 3 Tradeoff

Short-term

$2,500,000

$0

$1,666,667

Long-term

$1,000,000

$7,000,000

$3,000,000

To ensure that, along with spontaneous financing from accounts payable and accruals, adequate short-term financing will be available, Morton plans to establish an unsecured short-term borrowing arrangement with its local bank, Third National. The bank has offered either a line-of-credit agreement or a revolving credit agreement. Third National's terms for a line of credit are an interest rate of 2.50% above the prime rate, and the borrowing must be reduced to zero for a 30-day period during the year. On an equivalent revolving credit agreement, the interest rate would be 3% above prime with a commitment fee of 0.50% on the average unused balance.

Under both loans, a compensating balance equal to 20% of the amount borrowed would be required. The prime rate is currently 7%. Both the line-of-credit agreement and the revolving credit agreement would have borrowing limits of $1,000,000. For purposes of his analysis, Morton estimates that Kanton will borrow $600,000 on the average during the year, regardless of which financing strategy and loan arrangement it chooses. (Note: assume a 365-day year.)

TO DO

1.       Determine the total annual cost of each of the three possible financing strategies.

2.       Assuming that the firm expects its current assets to total $4 million throughout the year, determine the average amount of net working capital under each financing strategy. (Hint: Current liabilities equal average short-term financing.)

3.       Using the net working capital found in part b as a measure of risk, discuss the profitability-risk trade-off associated with each financing strategy. Which strategy would you recommend to Morton Mercado for Kanton Company? Why?

4.       Find the effective annual rate under: 1) the line-of-credit agreement and 2) the revolving credit agreement. (Hint: Find the ratio of the dollars that the firm will pay in interest and commitment fees to the dollars that the firm will effectively have use of.)

5.       If the firm expects to borrow an average of $600,000, which borrowing arrangement would you recommend to Kanton? Why?
Answer #1

1.Calculation of total annual cost Under each strategy : -

Aggressive Conservative Trade off
Short term@10% =$2500000*0.1=$250000 $0*0.1=$0 $1666667*0.1=$166667
Long term@14% =$1000000*0.14=$140000 =$7000000*0.14=$980000 $3000000*0.14=$420000
Total Cost $390000 $980000 $586667

2.Calculation of Net Working Capital : -

= Current Assets - Current liabilities

Aggressive Conservative Trade off
a)Current Assets $4000000 $4000000 $4000000
b)Current Liabilities $2500000 $0 $1666667
Net Working Capital(a)-(b) $1500000 $4000000 $2333333

3) Based on Networking capital Conservative strategy is Better strategy as it is Having high net working capital,as high net working capital is considered a sign of a well managed company with potential for growth.